A rectangle has a length that is 4 meters more than the width. The perimeter is 184 meters. Find the length
Answer by algebrahouse.com(1659)

x = width
x + 4 = length

Perimeter of a rectangle is 2(width) + 2(length)

2x + 2(x + 4) = 184 {substituted into perimeter formula}
2x + 2x + 8 = 184 {used distributive property}
4x + 8 = 184 {combined like terms}
4x = 176 {subtracted 8 from each side}
x = 44 {divided each side by 4}
x + 4 = 48 {substituted 44, in for x, into x + 4}

length is 48 m

For more help from me, visit: www.algebrahouse.com